Smoking by Race for Males Aged 18-24
             Smoker(S) Nonsmoker(N)  Row Total 
   White (W)  n(S&W)     n(N&W)        n(W) 
   Black (B)  n(S&B)     n(N&B)        n(B) 
Column Total  n(S)       n(N)        n(S or N) = n(W or B)  
	  
i.	P(S)	 	 
ii.	P(W)	 	 
iii.	P(S | W)	 	 
iv.	P(S | B)	 	 
v.	P(S and W)	 	 
vi.	P(N and B)


Answer by Edwin McCravy(20054)

             Smoker(S) Nonsmoker(N)  Row Total 
   White (W)  n(S&W)     n(N&W)        n(W) 
   Black (B)  n(S&B)     n(N&B)        n(B) 
Column Total  n(S)       n(N)        n(S or N) = n(W or B)  

----------------------------------------------------------

             Smoker(S) Nonsmoker(N)  Row Total 
   White (W)   278        555          833 
   Black (B)    11        156          167 
Column Total   289        711        1,000 

  i. P(S)      = 289/1000           = 0.289 
 ii. P(W)      = 833/1000           = 0.833
iii. P(S|W)    =  278/833           = 0.3337334934
iv. P(S|B)     =   11/167           = 0.0658682635   
v. P(S and W)  = 278/1000 = 139/500 = 0.278
vi. P(N and B) = 156/1000 =  39/250 = 0.156

Edwin